1. A line with an angle of inclination of 45° passes through (-5/2 , -9/2). If the ordinate of the point is 6, what is its abscissa?
  2. Find the angle from the line through (13/2, 0) and (2, -7) to the vertical line 5 units from the y-axis.
  3. The angle from line 1 to line 2 is arctan (2/3) and the slope of line 1 is -1. Find the slope of line 2.
  4. The inclination of line 1 is arctan (1/2). If line 2 makes an angle of 45° with line 1, find the slope of line 2.
  5. The vertices of a triangle are (7, 4), (1, 7) and (-3, -4) Find the interior angle at the vertex (-3, -4).




1
Expert's answer
2021-12-07T11:03:15-0500

1)

Gradient (m) of the line "=Tan45"

"=1"


Let "x=" abscissa

"\frac{\Delta\>Y}{\Delta\>x}=" gradient


"\frac{6-(-\frac{9}{2})}{x-(\frac{-5}{2})}=1"


"6+\frac{9}{2}=x+\frac{5}{2}"

"\implies\>x=8"



2)

Gradient of the line "=\frac{-7-0}{2-\frac{13}{2}}=\frac{7}{4.5}"


The angle that the line makes with horizontal

"=Tan^{-1}(\frac{7}{4.5})"

"=57.26"


Angle that the line makes with any vertical line

"=90-57.26\\=32.74"


3)

Angle between the lines

"=Tan^{-1}(\frac{2}{3})=33.69"


Angle between "L_1" and horizontal

"=Tan^{-1}(-1)\\=-45°"

Angle between "L_2" and horizontal

"=-45+" -"33.69"

"=-78.69"


Slope of "L_2=Tan(-78.69)"

"=-5"

Or between "L_2" and horizontal "=-45+33.69"

"=-11.31"


Slope of "L_2=" "Tan(-11.31)"

"=-0.2"


4) Angle between "L_2\>and\,L_1=Tan^{-1}(\frac{1}{2})"

"=26.565"

Angle between "L_2" and horizontal

"=45+26.565=71.565\\ Slope= tan(71.565)"

=3


5) side a"=\begin{vmatrix}
 1-7 \\
 7-4
\end{vmatrix}"


"\sqrt{(-6)^2+3^2}=√45"


Side b "=\begin{vmatrix}
 7--3\\
 4--4
\end{vmatrix}"


"=\sqrt{10^2+8^2)}=√164"


Side c="\begin{vmatrix}
 1--3\\
 7--4
\end{vmatrix}"


"=\sqrt{4^2+11^2}=√137"

Using cosine rule


"45=164+137-2\sqrt{164×137}\> cos\theta"


"Cos\>\theta=\frac{256}{299.8}"


"\theta=31.36"